Science Prep Night 20130102 UT


Closed for bad weather.

(Better formatted info on google drive)


Check and correct pointing, co-pointing using lbcrangebal
  • this will necessarily require running dofpia.

Run an OB (as in science mode - focus/collimate before hand) which uses guiding (i.e. exptime must be > 40 sec), dithering, and has stars on AO tech chip for tech-chip focusing to verify:
  • guiding
  • tech-chip focus corrections being sent
  • dithering is working as expected

Standard star fields if photometric (calib_OB) folders

Focal plane tilt confirmation good seeing? (not critical since we already have but MAY do)
  • superfoc
    • Obtain focus sequences for at least 2 position angles (0 and 180, e.g.)
    • OBs are in/home/LBTO/Calib_OBs/SUPERFOC/ are are: RB_rVsuperfoc.ob7, RB_rVsuperfoc7_pa90.ob,RB_rV_superfoc7_pa180.ob and RB_rV_superfoc7_pa270.ob.
    • Analyze in two ways (more details in TiltDataAnalysis)

run nmisc.starfoc for stars in subregions of each chip, usually top, middle, bottom, to produce an array of 12 best-focus values.

use script in LBTtools.Sandbox and gnuplot to generate and display this array of best-focus values.

(focfit3 uses Source Extractor to measure average FWHM in a bin. The number of bins along X and Y can be set).

4-chip extra/intra focal images

Obtain extra-/intra-focal image pairs for at least 2 position angles

OBs are in /home/LBTO/Calib_OBs/FOCUS and are:

RB_rVexin4.ob, RB_rVexin4_pa90.ob, RB_rVexin4_pa180.ob and RB_rVexin4_pa270.ob (extra then intra-focal image) and

RB_rVinex4.ob, RB_rVinex4_pa90.ob, RB_rVinex4_pa180.ob and RB_rVinex4_pa270.ob (takes intra- then extra-focal image)

Alternating exin and inex reduces the number of mirror focus (Z) moves and saves a few seconds.

More details on the TiltDataAnalysis


Closed dome: Organize/Purge MODS data disk must do this as user mods on mods1data, so ssh mods@mods1data periodically make UTdate subdirectories under /lhome/data and move data into these: cd /lhome/data mkdir UTdate (e.g. mkdir 20110929) mv mods1?.20110929.*.fits 20110929/ Total disk space on current system is about 100 Gb. A good nights data, with calibrations, will use 6-8 Gb. When the disk becomes full, you can delete older UTdate directories. du -sh (lists how much space is used by each subdirectory, in a human-readable form. To delete an entire directory rm -rf 20110616

I like to run, in addition to, a script which takes dual blue/red images through the imaging mask (execMODS /home/LBTO/SciPrep/MODS/ -- takes image of all masks

Open dome: Observe a spectrophotometric standard in all configurations -- alignment

Calibration scripts in ~MODSeng/modsScripts/modsSpecPhot If it is not photometric, try to remember to note this in the title (copy and edit the script) and ask the OSA to make sure to note it in their log. One of the primary spectrophotometric standards would be best, since these have better coverage in the red. This should take ~30 minutes - the stars have a range of brightnesses and, thus, exposure times (Note: we may make some versions to take only 1 image per configuration, so with nimgs 1). (Skipping for time and Target reasons; shouldn’t have changed) Check that offsets are accurate (scripts in /home/LBTO/SciPrep/MODS/AlignCheck/) 5-dice with detxy: execMODS dice5_xy_02+31.img execMODS dice5_relxy_02+31.img 5-dice with radec: execMODS dice5_relrd_02+31.img execMODS dice5_rd_02+31.img Blind offsets (script not prepared yet but dice5_relrd+02+31.img is a similar check). (For later) Measure Mt.Graham mean extinction coefficients for optical spectroscopy.

Preference for primary standards because of sampling at extreme wavelengths Need range of air masses (at least 1.25 - 2.20, our 25d limit is 2.367)


Closed dome: Check system time vs our GPS, at cmd line execute: "ntpdate -q" Check mask tables after exchange on the Mask Config GUI Make sure the column 'in use' does not have a check mark where there is no mask. Make sure there are no blank spaces before the mask name LMS files: Make sure the path name is correct Verify the lms files exist in this directory on the lucifer computer Shutdown and restart LUCIFER Control SW after mask exchange (already done) Take sieve mask images in N3.75 and N1.8 to check focus gradients Investigate possible loose lens in N1.80 camera Sieve mask, K filter Take image at 4 ROTANGLES for each EL Start with EL=90 and EL=25 (operating extremes) Inspect for any changes or strong asymmetries in IQ (top vs. bottom) Move calibration unit in: Check that lamps work Check gratings/tilts Take an imaging flat and check levels, dirt on window Move calibration unit out

On sky tests: Pupil image on sky Configure for this while changing instruments K band Scale exposure time so nothing saturates

MOS ACQUISITION problems (Dave -- might want to investigate here!)

Photometric standards (if photometric, of course) Sky brightness (J, H, K) photometric calibrated imaging, may need to add somewhat longer exposures to the standard star data

-- DavidThompson - 15 Jan 2013
Topic revision: r1 - 15 Jan 2013, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback